f Problem 11A (this is not in the textbook) Published reports indicate that a brain region call the nucleus accumbens (NAC) is involved in interval timing which is the perception of time in the seconds-to-minute range. To test this, researchers investigated whether removing the NAC interferes with rats' ability to time the presentation of a liquid reward. Using a conditioning procedure, the researchers had rats press a lever for reward that was delivered after 16 seconds. The time that eight rats responded the most (peak responding) was recorded before and after a surgery to remove the NAC. The peak responding times are given in the following table: Peak Interval Timing Before NAC Surgery After NAC Surgery 15 14 16 16 17 18 15 16 20 26 20 18 21 18 23 Does the data show that there is a significant difference in peak interval timing before and after NAC surgery? Complete the four steps and use a-05 to complete. Compute eta squared to measure the size of the treatment effect. Write a sentence demonstrating how the results of the hypothesis test and the measure of effect size would appear in a research report. a. b. c. SPSS Problems Instructions on how to review SPSS and answer the question are at the end of each chapter. 1. Two biology teachers are in competition with each other to see which type of class is better for students to learn in. Instructor A teaches Biology 101 as an online course 3 days a week. Instructor B also teaches Biology 101 3 days a week but his class is in person. Do final grades differ between Instructor A and Instructor B? Group Statistics instructor N Mean Std. Deviation Std. Error Mean final A 7 79.5714 7.955233.00680 2.08003
Independent Samples Test Levene's Test for Equality of Variances ttest for Equality of Means 95% Confidence Interval Sig, 02- Mean St. Error ofthe Difference df tallen Difference Difference- Lower-_Jpper final Equal 2.586 .134 12 004 -12.85714 3.65614-20.82318 4.89110 vanances assumed Equal variances not assumed 3.517 -10.673012.85714 3.65614 -20.93447 4.77981 3.517 1. Based on the value of the test statistic, state the decision for a two-independent-sample t test. 2. State the conclusion for this test in APA format. Make sure you summarize the test statistic and the effect size. Provide an interpretation for the statistics for your report. See page 316-319 on how this is written.
Answer #1

using equal variance

(a) here we use t-test with

null hypothesis H0:mean1=mean2 and alternate hypothesis H1:mean1≠mean2

statistic t=|(mean1-mean2)|/SE(difference) =3.517 with df is n=n1+n2-2 =12and sp2=((n1-1)s12+(n2-1)s22)/n

since the two-tailed p-value=0.004 ( for unequal varaince p-value=0.005) is less than typical alpha=0.05, so we reject H0 and conclude that there is significant difference between before and after

(b) Eta squared =t2/(t2+df)= t2/(t2 + (n1+n2-2))=(3.517*3.517)/(3.517*3.517+14)=0.4691

(c) since eta square is large so there is significant difference between before and after

  • .02 ~ small
  • .13 ~ medium
  • .26 ~ large